# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

### v2

@@ I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
@@
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a class of hybrid systems that involve state-triggered switching. However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS). MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16]. [17] is an excellent reference on the classical results for discrete-time MJLS.
Constrained control of MJLS has been an active area of research, e.g. with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20]. Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
@@ II MJLS Dynamics
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\mathcal{X}_{k}(i)\triangleq\mathbb{E}\left[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}\right],$
$X_{k}\triangleq\mathbb{E}[x_{k}x_{k}^{\top}]=\sum_{i\in\Omega}\mathcal{X}_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
We assume that the initial probability distribution $\rho_{0}\in\mathbb{R}^{1\times N}$ is given. Therefore, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$.
Remark 1.
Note that $x_{k}$ is independent of $r_{k}$. Therefore, $q_{k}(i)=\mathbb{E}[x_{k}]\mathbb{E}[\mathds{1}_{r_{k}=i}]=\mu_{k}\rho_{k}(i)$, and $\mathcal{X}_{k}(i)=\rho_{k}(i)X_{k}$.
We assume a control policy of the following form:
$u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k})$ (2)
@@
$\rho_{k}(i)>0,\;\forall k\in\mathbb{Z}_{0}^{T},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
@@
The covariance of the given MJLS is positive definite at all time steps, i.e. $\Sigma_{k}\succ 0$ $\forall k\in\mathbb{Z}_{0}^{T}$.
II-A Mean Propagation
We now focus on MJLS mean dynamics. We know:
$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
$=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
@@ II-A Mean Propagation
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0}^{T-1}}$ which satisfy eq. 4.
II-B Covariance Propagation
For covariance propagation, we first consider $\mathcal{X}_{k+1}(j)$:
$\mathcal{X}_{k+1}(j)=\mathbb{E}\left[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}\right]$ (5)
$=\sum_{i\in\Omega}\left\{p_{ij}\mathbb{E}\left[\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)\right.\right.$
$\left.\left.\left(A_{k}(i)x_{k}+B_{k}(i)u_{k}(i)+c_{k}+G_{k}(i)w_{k}\right)^{\top}\mathds{1}_{r_{k}=i}\right]\right\}$
Using identities derived in the appendix, we can simplify eq. 5, and substitute that in $\Sigma_{k+1}=\sum_{j\in\Omega}\mathcal{X}_{k+1}(j)-\mu_{k+1}\mu_{k+1}^{\top}$ to get:
@@ II-B Covariance Propagation
Separately analyzing some terms inside the summation in eq. 6:
$A_{K}(i)\mu_{k}\mu_{k}^{\top}A_{K}^{\top}(i)+A_{k}(i)\mu_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+A_{k}(i)\mu_{k}c_{k}^{\top}$
$+B_{k}(i)\overline{u}_{k}(i)\mu^{\top}_{k}A_{k}^{\top}(i)+B_{k}(i)\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)$
$+B_{k}(i)\overline{u}_{k}(i)c^{\top}_{k}+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}B_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
$=(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})(A_{k}(i)\mu_{k}+B_{k}\overline{u}_{k}(i)+c_{k})^{\top}$
Note these terms are linear in $p_{ij}$ while $\mu_{k+1}\mu_{k+1}^{\top}$ involves the product $p_{ij}p_{ij}$, explaining why they don’t cancel out. This demonstrates why the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
@@ III-A Optimal Control Problem
$\min$ $\;J=\sum_{k=0}^{T-1}J_{k}$ (7)
$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0}^{T-1},$
$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+G_{k}(r_{k})c_{k}$
$x_{0}$ $\sim\mathcal{N}(\overline{\mu}_{0},\overline{\Sigma}_{0})$
$\mathbb{E}[x_{T}]=$ $\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$
where $J_{k}=\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ with $Q_{k}\succeq 0$ and $R_{k}\succ 0$, $\forall k\in\mathbb{Z}_{0}^{T-1}$. Simplifying $J_{k}$, we get:
$J_{k}$ $=\sum_{i\in\Omega}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
$=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
@@ III-B Convex Relaxation
The nonlinear terms in eq. 6 result in a non-convex optimization landscape for eq. 7. To solve it in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 6:
@@ III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 1 and then uses its solution to solve Problem 2.
Problem 1.
@@
$H_{K}^{\Sigma}=\Sigma_{k+1}$ $-F_{\tau}(\Xi)=0$ (17b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (17c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (17d)
@@
Slater’s condition is satisfied for Problem 2, i.e. there exists a strictly feasible solution for Problem 2.
@@
Under Assumptions 1, 2 and 3, the optimal solution to Problem 2 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
@@
For Problem 2, we can define the Lagrangian as:
@@
where $C_{k}(i)$ is defined in eq. 12, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 17b which shares the same symmetry properties with eqs. 11 and III-B. $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
@@
If we attempt to carry out similar relaxations for the mean quadratic terms to relax the covariance propagation equation (6) without assuming the mean trajectory $\tau$ is given, we find that it is not possible to prove that these relaxations are lossless. In fact, we get a condition where $M_{k}^{(3)}=-M^{(4)}_{k}$, where $M_{k}^{(3)}$ and $M^{(4)}_{k}$ are Lagrange multipliers for respective relaxations. Since they are both positive semi-definite, it implies $M_{k}^{(3)}=M^{(4)}_{k}=0$. Thus, we cannot use our existing arguments to prove the relaxation is lossless. Numerical simulations also demonstrated this lossy behavior, motivating our two-step sequential mean-covariance optimization approach.
@@ IV Chance-Constrained MJLS Covariance Steering
Previous works [23, 8] have focused on deterministic and convex reformulations of eqs. 20 and 21, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [24] between the MJLS distribution and the approximating Gaussian. It has been shown previously [25] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 20 and 21.
@@ IV-A Deterministic Formulation
Note that while eqs. 22 and 23 are convex in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
@@ IV-B Iterative Convex Control Synthesis Framework
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i),\zeta_{m}}J^{c}_{\Sigma}\hskip 56.9055pt$ (27a)
@@
$H_{K}^{\Sigma}=\Sigma_{k+1}-F_{\tau}(\Xi)$ $=0$ (27b)
$\begin{bmatrix}Y_{k}(i)&L_{k}(i)\\ L_{k}^{\top}(i)&{\Sigma}_{k}\end{bmatrix}$ $\succeq 0$ (27c)
$\Sigma_{0}=\overline{\Sigma}_{0},$ $\quad\Sigma_{T}\preceq\Sigma_{f}$ (27d)
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $\leq\zeta_{1,j},$ (27e)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $\leq\zeta_{2,j},$ (27f)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $\leq\zeta_{3,j}$ (27g)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $\leq\zeta_{4,j}$ (27h)
@@
Slater’s condition is satisfied in Problem 4 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 27c remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 4.
@@
Under Assumptions 1, 2 and 4, the optimal solution to Problem 4 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
@@
We can define the Lagrangian for Problem 4 as follows:
@@
where $C_{k}(i)$ is given in eq. 12 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 27e, 27f, 27g and 27h respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 29c. Rearranging eq. 29a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 9, we simplify eq. 29b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the left hand side as $\Delta_{k}(i)$ and taking the determinant of both sides, we get:

### v3

@@ I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon problem and has attracted significant interest recently[2, 3, 4, 5].
@@
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a subclass of hybrid systems but don’t extend to Markov jump linear systems (MJLS). MJLS is a type of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature[14] with applications in fault-tolerant control [15], networks [16], and macroeconomic models [17].
Constrained control of MJLS has been an active area of research, and has benefited from the enormous progress in constrained control of single mode systems [bemporad_explicit_2002, 18] e.g. with state and control polytopic constraints for linear time-invariant[19] and periodic systems[20], and with second-moment constraints[21]. Chance constraints have previously been incorporated for MJLS[22], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
@@ II Problem Statement
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$. For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$. We approximate $q_{k}(i)$ by $\rho_{k}(i)\mu_{k}$, which was found to be a reasonable approximation numerically. Similarly, we approximate $\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]$ by $\rho_{k}(i)\Sigma_{k}$. The appendix section elaborates on this approximation and the consequent justification. We now define the MJLS covariance steering optimal control problem below:
Problem 1.
Solve:
$\min_{x_{k},u_{k}(i)}$ $\;J=\sum_{k=0}^{T-1}\mathbb{E}\left[x_{k}Q_{k}x_{k}^{\top}+u_{k}(i)R_{k}u_{k}^{\top}(i)\right]$ (2a)
$s.t$$.\;\;\;\forall k\in\mathbb{Z}_{0:T-1},$
$x_{k+1}$ $=A_{k}(r_{k})x_{k}+B_{k}(r_{k})u_{k}+c_{k}+G_{k}(r_{k})w_{k}$ (2b)
$x_{0}\sim\mathcal{N}(\overline{\mu}_{0},$ $\overline{\Sigma}_{0}),\;\mathbb{E}[x_{T}]=\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$ (2c)
$P(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad P(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\Omega$. The distribution $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
@@
$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
@@
$\Sigma_{k}\succ 0$ $\forall k\in{\mathbb{Z}_{0:T}}$.
Problem 1 is hard to solve because of stochastic decision variables. To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS. We consider a mode-dependent affine feedback control policy:
${u_{k}(i)=\overline{u}_{k}(i)+K_{k}(i)\left(x_{k}-\overline{x}_{k}(i)\right)}$ (3)
Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously. Focusing on the mean propagation first, we consider $q_{k}(i)$ below:
$q_{k+1}$ $(j)=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
@@
Similarly, covariance propagation is studied below:
$\Sigma_{k+1}$ $=\left(\sum_{i\in\Omega}\mathbb{E}[(x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k}=i}]\right)-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
Substituting the identities given in the appendix, we get:
@@
eq. 7 shows that the mean-covariance separation of single-mode dynamics ([10]) does not extend to MJLS.
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in{\mathbb{Z}_{0:T-1}}}$ which satisfy eq. 5. We now define our optimal control problem in terms of $\mu_{k}$ and $\Sigma_{k}$, which approximates Problem 1.
Problem 2.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (8)
$s.t.\;\forall k\in\mathbb{Z}_{0:T-1},$
[…]
We write $J_{k}$ in terms of $\mu_{k}$ and $\Sigma_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
The nonlinearity in eq. 7 make Problem 2 results in a non-convex optimization landscape, while the chance constraints eq. 2d are not yet represented in closed form. We tackle these challenges in the following sections by first proposing a tight convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain tight with a deterministic formulation of chance constraints too.
@@ III-A Optimal Control Problem
Problem 3.
Solve:
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (10)
[…]
$\;\;\mu_{0}=\overline{\mu}_{0}\;\;\mu_{T}=\mu_{f}$ (11)
@@ III-B Convex Relaxation
To solve Problem 3 in a computationally tractable manner, we propose a two-step optimization process, which optimizes the mean and covariance variables sequentially using convex relaxations. We first introduce the following change of variables to convexify eq. 7:
@@ III-C Two-step Convex Optimization Framework
We now define the two subproblems which are solved sequentially. Therefore, the two-step optimization process first solves Problem 4 and then uses its solution to solve Problem 5.
Problem 4.
@@
$\Sigma_{0}=\overline{\Sigma}_{0},\quad\Sigma_{T}\preceq\Sigma_{f}$ (20c)
@@
Slater’s condition is satisfied for Problem 5, i.e. there exists a strictly feasible solution for Problem 5.
@@
Under Assumptions 1, 2 and 3, the optimal solution to Problem 5 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$, hence, lossless.
@@
For Problem 5, we can define the Lagrangian as:
@@
where $C_{k}(i)$ is defined in eq. 15, $\Lambda_{k}$, $M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively. $\Lambda_{k}$ is symmetric because of symmetry in eq. 14 and $M_{k}$ and $\lambda_{F}$ are symmetric by definition. The relevant first-order KKT conditions for optimality are:
@@
Similar relaxations cannot be performed for mean quadratic terms in the eq. 7 because they involve rank-1 matrices which doesn’t allow the use of Schur’s complement to relax the non-convex substitutions as LMIs, motivating our two-step sequential mean-covariance optimization approach.
@@ IV Chance-Constrained MJLS Covariance Steering
Previous works [24, 8] have focused on deterministic and convex reformulations of eqs. 23 and 24, but they usually require Gaussian dynamics. MJLS dynamics result in a mixture of exponential Gaussians ($N^{T}$ Gaussians for $N$ modal MJLS at time $T$.) In this work, we approximate the MJLS state by a single Gaussian distribution. To achieve this, we seek to minimize the Kullback-Leibler divergence [25] between the MJLS distribution and the approximating Gaussian. It has been shown previously [26] that this is achieved by performing moment matching, i.e. setting the first two moments of the approximating Gaussian to be the same as the original distribution which is being approximated. We leverage this useful property to approximate the state and control of MJLS using Gaussian distributions. We approximate $x_{k}$ by $\mathcal{N}(\mu_{k},\Sigma_{k})$, using moment matching to the approximating Gaussian. Similarly, for control in mode $i$, $\mathbb{E}[u_{k}(i)]=\overline{u}_{k}(i)$ and $\mathbb{E}[(u_{k}(i)-\overline{u}_{k}(i))(u_{k}(i)-\overline{u}_{k}(i))^{\top}]=K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)=Y_{k}(i)$, therefore $u_{k}(i)$ can be approximated by $\mathcal{N}(\overline{u}_{k}(i),Y_{k}(i))$. We now discuss the deterministic formulation of eqs. 23 and 24.
@@ IV-A Deterministic Formulation
Note that while eqs. 25 and 26 are linear in the mean variables, they are nonconvex in the covariance terms. We convexify them by rearranging and squaring both sides as:
@@ IV-B Iterative Convex Control Synthesis Framework
$\min_{{\Sigma}_{k},L_{k}(i),Y_{k}(i)}J^{c}_{\Sigma}\hskip 56.9055pt$ (30a)
@@
[…]
$m^{2}_{\mathcal{N}}(\delta_{x,j}){a_{j}^{\top}\Sigma_{k}a_{j}}-(a_{j}^{\top}\mu_{k}+b_{j})^{2}$ $-\zeta_{1,j}\leq 0,$ (30c)
$m^{2}_{\mathcal{N}}(\delta_{u,j}){p_{j}^{\top}Y_{k}(i)p_{j}}-(p_{j}^{\top}\mu_{k}+d_{j})^{2}$ $-\zeta_{2,j}\leq 0,$ (30d)
$m^{2}_{\chi^{2}}(\epsilon_{x},n_{x})\lambda_{\max}(\Sigma_{k})$ $-\zeta_{3,j}\leq 0$ (30e)
$m^{2}_{\chi^{2}}(\epsilon_{u}(i),n_{u})\lambda_{\max}(Y_{k}(i))-u_{m}^{2}$ $-\zeta_{4,j}\leq 0$ (30f)
@@
Slater’s condition is satisfied in Problem 7 for $\zeta_{m}=0$. This ensures the convex relaxation in eq. 16 remains lossless while satisfying chance constraints at $\zeta_{m}=0$.
We now show that the convex relaxation $Y_{k}(i)\succeq L_{k}(i)\Sigma_{k}^{-1}L_{k}^{\top}(i)$ remains lossless in Problem 7.
@@
Under Assumptions 1, 2 and 4, the optimal solution to Problem 7 satisfies $Y_{k}(i)=L_{k}(i){\Sigma}^{-1}_{k}L_{k}^{\top}(i)$.
@@
We can define the Lagrangian for Problem 7 as follows:
@@
where $C_{k}(i)$ is given in eq. 15 and $e_{x,h}$, $e_{u,h}$, $e_{x,n}$, and $e_{u,n}$ represent eqs. 30c, 30d, 30e and 30f respectively with $\lambda_{x,h}$, $\lambda_{u,h}$, $\lambda_{x,n}$, and $\lambda_{u,n}$ being the respective Lagrange multipliers. $\Lambda_{k}$,$M_{k}$ and $\lambda_{F}$ are Lagrange multipliers corresponding to $H_{k}^{\Sigma}=0$, $C_{k}(i)\preceq 0$ and $\Sigma_{T}\preceq\Sigma_{f}$ respectively, and are all symmetric as explained in the proof of theorem 1. The relevant first-order KKT conditions for optimality are:
@@
where $\pi^{(k)}_{ij}=p_{ij}\rho_{k}(i)$ and $v_{k}(i)$ is the eigenvector of $Y_{k}(i)$ corresponding to its maximum eigenvalue. We now prove by contradiction that $C_{k}(i)=0$, or the relaxation is lossless. Let us suppose that $C_{k}(i)$ has at least one nonzero eigenvalue, which implies $M_{k}(i)$ is singular from Lemma 1 and eq. 32c. Rearranging eq. 32a, we get $\sum_{j\in\Omega}\pi^{(k)}_{ij}B_{k}^{\top}(i)\Lambda_{k}=-M_{k}(i)L_{k}(i){\Sigma}_{k}^{-1}A_{k}^{-1}(i)$. Using this and the substitutions defined in eq. 12, we simplify eq. 32b as: $\rho_{k}(i)R_{k}+\lambda_{u,n}m^{2}_{\chi^{2}}(\epsilon_{u},n_{u})v_{k}(i)v_{k}^{\top}(i)+\lambda_{u,h}m^{2}_{\mathcal{N}}(\delta_{u,j})p_{j}p^{\top}_{j}=M_{k}(i)(I+K_{k}(i)A_{k}^{-1}(i)B_{k}^{\top}(i))$. Representing the LHS as $\Delta_{k}(i)$ and taking the determinant of both sides, we get: